# Towards Neural Scaling Laws for Time Series Foundation Models

### v1

@@ C.1 Training Setup
Based on the constructed datasets of different sizes, we trained a series of encoder-only Transformers and decoder-only Transformers with increasing data sizes and model parameters. To improve batch processing efficiency and handle varying sequence lengths, we employ sequence packing (Raffel et al., 2020), which reduces padding requirements. Moreover, we train models with forecasting capabilities over varying context and prediction lengths. Rather than defining a fixed context and prediction length, we sample 15% - 50% lengths as forecast horizon and the remaining as context horizon, for a given time series. As most probabilistic forecasting models, we minimize the negative log-likelihood of the predicted patch with respect to the ground truth.
@@ C.2 Evaluation Setup
We primarily used the LSF (Wu et al., 2023) and Monash (Godahewa et al., 2021) datasets to evaluate the model’s out-of-distribution generalization ability. For the LSF dataset, we only used the test set and performed evaluations in a non-overlapping rolling window fashion, with the stride equal to the prediction length. For the Monash dataset, considering its large size, we set only one window per sample for evaluation. The evaluated results in Monash dataset are exhibited in Appendix D.3. The below table details the evaluation setting per dataset:
@@
Dataset Domain Frequency # Prediction Length # Rolling Evaluations.
@@
Traffic Weekly Transport W 8 862
Australian Electricity Demand Energy 30T 336 5
@@ C.3 Evaluation Metrics
Mean Absolute Percentage Error (MAPE). MAPE is a commonly used metric to evaluate the accuracy of a forecasting model by expressing the error in percentage terms, making it unit-free and easily interpretable. It measures the average magnitude of the absolute percentage error between the predicted and actual values, relative to the actual values. For a univariate time series, the error is defined as:
@@
Figure 9: Training encoder-only models show that, as model scale increases, larger models require fewer samples to achieve the same level of performance.
Figure 10: Scaling Pattern Depends on Data Distributions. The blue, red, and green plots illustrate the scaling behavior of log-likelihood loss on in-distribution (ID) data, out-of-distribution (OOD) LSF data, and OOD Monash data, respectively. We observe that the model exhibits different scaling patterns on the two different OOD datasets, indicating that OOD scaling behaviors significantly depend on the relationship between the unseen data distribution and the training data distribution.
@@ D.1 Emergent behaviors
Fig. 8 presents case studies on ”emergent abilities” of TSFMs for out-of-distribution (OOD) predictions. As model size increases, one would expect the MAPE to follow a smooth power-law decline, indicating continuous improvement. However, the plots reveal distinct phenomena, where performance changes abruptly rather than gradually. This behavior is akin to the “emergent behaviors” described in (Wei et al., 2022), where certain capabilities absent in smaller models only appear in larger ones. The emergence of improved performance at certain model sizes could indicate the development of higher-order patterns or representations within the model. These emergent abilities might stem from the model’s capacity to recognize and generalize complex temporal patterns that were previously inaccessible at smaller scales. Such abilities is evident particularly in zero-shot OOD predictions, where the model must extrapolate from training data to entirely unseen scenarios.
@@ D.2 Sample Efficiency
Fig. 9 illustrates the evaluation results of different-sized encoder-only models on a validation set during training. The key observation is that larger models achieve better performance with fewer training steps compared to smaller models, demonstrating higher sample efficiency. This is important in scenarios where data are limited, as larger models can achieve superior performance without processing as much data.
@@ D.3 Scaling pattern depends on data distributions
Fig. 10 highlights the influence of different data distributions on the scaling behavior of the model. By comparing the green line, representing the Monash dataset, and the red line, representing the LSF dataset, we observe a relatively consistent offset and similar slopes across the scaling curves. This suggests that, when transferring the model from the training data distribution to other OOD distributions, there is a predictable decrease in performance that varies by the dataset. However, the model’s gains from scaling—whether by increasing model size, compute, or dataset size—follow a fixed proportional relationship. This indicates that while OOD performance degrades, scaling the model still yields a consistent improvement ratio across different distributions, albeit with varying levels of absolute performance. Moreover, in the future, we can further analyze the reasons why transfer cost arises and establish a law to predict the transfer performance loss in population-risk based theoretical analysis (Yang et al., 2021) through Wasserstein measurement.
D.4 Scaling pattern depends on performance metrics
@@
We investigate the scaling behaviors of five common performance metrics: log-likelihood, MAPE, CRPS, sMAPE and MAE(as shown in Figures 11 12 and 13). All metrics exhibit a decreasing trend following an approximate power-law; however, each metric demonstrates distinct scaling characteristics, reflected in their varying power-law exponents. Previous work by Ghorbani et al. (2021) establishes a relationship between large language models’ log-likelihood loss and the BLEU score in translation tasks. Similarly, future research could explore a transformation between log-likelihood loss and time series forecasting metrics, offering a means to predict forecasting performance from training loss.
Figure 11: Scaling Pattern of CRPS. This figure illustrates scaling laws for CRPS across three training factors: model size, compute, and dataset size. The blue plots indicate the ID performance, while the red plots represent OOD performance. In both ID and OOD scenarios, CRPS decreases with the three factors scaling up.
Figure 12: Scaling Pattern of sMAPE. This figure presents the scaling laws for sMAPE in relation to three training factors: model size, compute, and dataset size. The blue plots show performance in ID scenarios, while the red plots depict OOD performance. In both cases, sMAPE decreases as all three factors increase.
Figure 13: Scaling Pattern of MAE. This figure illustrates scaling laws for MAE in relation to model size, compute, and dataset size. The blue lines represent ID performance, while the red lines show OOD performance. In all cases, increasing model size, compute, or dataset size reduces MAE, with out-of-distribution performance improving more significantly.

### v2

@@ C.1 Training Setup
Based on the constructed datasets of different sizes, we trained a series of encoder-only Transformers and decoder-only Transformers with increasing data sizes and model parameters. To improve batch processing efficiency and handle varying sequence lengths, we employ sequence packing (Raffel et al., 2020), which reduces padding requirements. Moreover, we train models with forecasting capabilities over varying context and prediction lengths. Rather than defining a fixed context and prediction length, we sample 15% - 50% lengths as forecast horizon and the remaining as context horizon, for a given time series. As most probabilistic forecasting models, we minimize the negative log-likelihood of the predicted patch with respect to the ground truth. The loss convergence process of encoder-only models and decoder-only models is shown in Figure 10. As the number of parameters increases, both encoder-only and decoder-only models tend to converge faster to a lower training loss.
Figure 10: Convergence Process of Training Loss. This figure show the training loss convergence for encoder-only (left) and decoder-only (right) models of various sizes, ranging from 1K to 1B parameters. Larger models demonstrate faster convergence and lower final training loss, highlighting their improved capacity for optimization.
@@ C.2 Evaluation Setup
We primarily used the LSF (Wu et al., 2023) and Monash (Godahewa et al., 2021) datasets to evaluate out-of-distribution generalization ability. For the LSF dataset, we only used the test set and performed evaluations in a non-overlapping rolling window fashion, with the stride equal to the prediction length. For the Monash dataset, considering its large size, we set only one window per sample for evaluation. Table LABEL:tab:ood_dataset details the evaluation setup per dataset.
Following the statistical method in (Qiu et al., 2024), we compute several statistical characteristics of the datasets used in this study. Table 4 presents the mean statistical features of the pre-training, LSF, and Monash datasets. SNR refers to the signal-to-noise ratio, which quantifies the proportion of meaningful signal relative to random noise in a time series. Shifting captures changes in the probability distribution of a time series over time, with values closer to 1 indicating more severe distributional shifts. Stationarity describes a property of a time series where the mean remains constant over time, and the variance is finite and stable across observations. Transition reflects the regular and identifiable fixed features within a time series, such as trends, periodicity, or other predictable patterns. Overall, our pre-training data and OOD test data both show a high quality and stationarity.
@@
Dataset Domain Frequency # Prediction Length # Samples.
@@
Table 4: Mean Statistic Characteristics.
Dataset SNR Shifting Stationarity Transition
Pre-training 25.47 0.33 0.47 0.03
LSF 19.78 0.15 0.50 0.01
Monash 19.84 0.34 0.38 0.09
@@ C.3 Evaluation Metrics
Mean Absolute Percentage Error (MAPE). MAPE is a commonly used metric to evaluate the accuracy of a forecasting model by expressing the error in percentage terms, making it unit-free and easily interpretable. It measures the average magnitude of the absolute percentage error between the predicted and actual values, relative to the actual values. For a uni-variate time series, the error is defined as:
@@
Mean Absolute Scaled Error (MASE). MASE is a scale-independent metric used to evaluate forecasting accuracy by comparing the absolute error of a forecast to the average error of a naive forecasting method, such as a one-step-ahead random walk. MASE provides a robust way to assess model performance across different datasets with varying scales, addressing the limitations of metrics like MAPE that are sensitive to the magnitude of actual values. For a uni-variate time series, the MASE of the $i$-th time series is calculated as:
$\text{MASE}=\frac{\frac{1}{H}\sum_{j=t+1}^{t+H}\lvert e_{j}^{(i)}\rvert}{\frac{1}{T-1}\sum_{k=2}^{T}\lvert y_{k}^{(i)}-y_{k-1}^{(i)}\rvert}$
where $|y_{k}^{(i)}-y_{k-1}^{(i)}|$ represents the difference between consecutive values in the naive forecast. MASE is particularly useful because it accounts for the variability and scale of the time series, making it comparable across datasets. Unlike MAPE, MASE does not produce extreme values for small or zero actual values, making it more stable and reliable for time series evaluation.
@@ Appendix D Additional Results
Figure 11: Emergent Behaviors in us birth Dataset. The sub-figures, arranged from left to right and top to bottom, show prediction results for models of sizes 10K, 1M, 10M, and 100M. Models with sizes 10K and 1M are unable to capture the periodic pattern, whereas models with sizes 10M and 100M accurately reflect it.
Figure 12: Emergent Behaviors in nn5 daily Dataset. From left to right and top to bottom, the sub-figures illustrate the prediction outcomes of models with sizes 10K, 1M, 10M, and 100M, respectively. While the smaller models (10K and 1M) fail to recognize the periodic pattern in the data, the larger models (10M and 100M) effectively capture and reflect the periodic structure.
Figure 13: Emergent Behaviors in australian electricity Dataset. From left to right and top to bottom, the sub-figures display the prediction results of models with sizes 10K, 1M, 10M, and 100M. The models with sizes 10K and 1M fail to accurately capture the periodic pattern, while those with sizes 10M and 100M successfully do so.
Figure 14: The figure illustrates the validation NLL during training for encoder-only models of different sizes. Larger models tend to achieve lower NLL with processing the same time points.
Figure 15: Scaling Pattern of NLL. This figure illustrates scaling laws for NLL in relation to model size, compute, and dataset size. The blue lines represent ID performance, while the red and green lines show OOD performance on LSF subset and Monash subset.
Figure 16: Scaling Pattern of MAPE. This figure presents scaling laws for MAPE as functions of model size, compute, and dataset size. ID performance is illustrated by the blue lines, while OOD performance is depicted in red for the LSF subset and green for the Monash subset.
Figure 17: Scaling Pattern of SMAPE. This figure depicts the scaling laws for SMAPE with respect to model size, computational resources, and dataset size. ID performance is marked by blue lines, while red and green lines correspond to OOD performance on the LSF and Monash subsets, respectively.
Figure 18: Scaling Pattern of MASE. The figure shows how MASE scales with model size, compute, and dataset size. ID performance is indicated by blue lines, whereas OOD performance is highlighted in red and green for the LSF and Monash subsets.
Figure 19: Scaling Pattern of CRPS. The figure illustrates the scaling laws for CRPS in relation to model size, compute, and dataset size. Blue lines represent ID performance, while red and green lines depict OOD performance on the LSF and Monash subsets, respectively.
@@ D.1 Emergent behaviors
Figure 8 presents case studies on ”emergent abilities” of TSFMs for OOD predictions. As model size increases, one would expect the MAPE to follow a smooth power-law decline, indicating continuous improvement. However, the plots reveal distinct phenomena, where performance changes abruptly rather than gradually. Figures 11 - 13 visually show the transition from the small model prediction results to the large model prediction results. We see that the 10K and 1M models are unable to capture the historical seasonal patterns to make accurate predictions; however, when the model size reaches 10M, it can make accurate predictions, far better than the small model. This behavior is akin to the “emergent behaviors” described in (Wei et al., 2022), where certain capabilities absent in smaller models only appear in larger ones. The emergence of improved performance at certain model sizes could indicate the development of higher-order patterns or representations within the model. These emergent abilities might stem from the model’s capacity to recognize and generalize complex temporal patterns that were previously inaccessible at smaller scales. Such abilities is evident particularly in zero-shot OOD prediction, where the model must extrapolate from training data to entirely unseen scenarios.
@@ D.2 Sample Efficiency
Figure 14 illustrates the evaluation results on the ID data for encoder-only models of different sizes during training. The key observation is that larger models achieve better performance with fewer training steps compared to smaller models, demonstrating higher sample efficiency. This is important in scenarios where data are limited, as larger models can achieve superior performance without needing to process as much data. An explanatory intuition is that, when the number of parameters is large, even a small change in parameters can cause a large change in outputs. From a theoretical perspective, the Neural Tangent Kernel (Jacot et al., 2018) framework provides insights into why this occurs. As neural networks become infinitely wide, their behavior approximates that of a kernel method, enabling them to effectively minimize the data distribution’s generalization error. This allows large models to generalize well with fewer samples, further supporting their observed sample efficiency in empirical settings. Moreover, in NLP (Kaplan et al., 2020) and CV (Zhai et al., 2022) fields, larger models have demonstrated higher sample efficiency than smaller models.
@@ D.3 Scaling pattern depends on data distributions
In Figures 15 - 19, we present the scaling laws evaluated on the Monash subset, along with results from the LSF subset and the ID test data. By comparing the green line, representing the Monash dataset, and the red line, representing the LSF dataset, we observe a relatively consistent offset and similar slopes across various metrics. This suggests that, when transferring the model from the training data distribution to other OOD distributions, there is a predictable decrease in performance that varies by the target dataset. However, the model’s gains from scaling—whether by increasing model size, compute, or dataset size—follow a fixed proportional relationship. This indicates that while OOD performance degrades, scaling the model still yields a consistent improvement ratio across different distributions, albeit with varying levels of absolute performance. Moreover, in the future, we can further analyze the reasons why transfer cost arises and establish a law to predict the transfer performance loss in population-risk based theoretical analysis (Yang et al., 2021) through Wasserstein measurement.
D.4 Scaling pattern depends on performance metrics.
@@
We investigate the scaling behaviors of five common performance metrics: NLL, MAPE, SMAPE, MASE, and CRPS, (as shown in Figures 15 - 19). All metrics exhibit a decreasing trend following an approximate power-law; however, each metric demonstrates distinct scaling characteristics, reflected in their varying power-law exponents. Previous work by (Ghorbani et al., 2021) establishes a relationship between large language models’ log-likelihood loss and the BLEU score in translation tasks. Similarly, future research could explore a transformation between log-likelihood loss and time series forecasting metrics, offering a means to predict forecasting performance from training loss.
Appendix E Further Discussion
Scaling Laws for Multivariate Time Series Forecasting. Extending our findings to multivariate time series forecasting is promising but presents some challenges. First, large and diverse multivariate datasets are limited, unlike the more standardized uni-variate time series datasets. This scarcity makes it difficult to conduct experiments that scale both model size and data volume in a controlled manner. Second, establishing scaling laws for multivariate time series require analyzing the impact of variable count and correlation strength, increasing the experimental complexity. Third, multivariate foundation models lag behind uni-variate ones, facing challenges in designing architectures that accommodate variable counts and capture complex inter-variable dependencies.
Pre-training Data Mixing Strategy. It’s significant to investigate the impact of data mixing strategies on model performance, as different strategies can introduce performance bias or even degrade model performance. Time series pre-training datasets, often comprising data from dozens to hundreds of sources, create exponentially large possibilities for combinations. Instead of exhaustive empirical research, a more practical approach is to establish an extrapolable analytical framework that balances experimental complexity (Hashimoto, 2021; Ye et al., 2024). Additionally, deeper analysis of time series dataset characteristics is essential. This would provide insights into the nature of time series data and facilitate the development of better data organization formats, moving beyond simple combination strategies.
Impact of Model Architecture on Scalability. Our research comparing the scalability of encoder-only and decoder-only Transformers can serve as the groundwork for future exploration of other architectures. CNN-based models (Wu et al., 2022), RNN-based models (Gu et al., 2021), and hybrid models (Lieber et al., 2024) each demonstrate unique strengths in sequence modeling. Investigating their scalability within the context of time series modeling is a promising direction for future research, and we hope our findings will inspire further work in this area.
Impact of Specific Module on Scalability. Our case studies on the scalability of Chronos and Moirai demonstrate that certain module designs in these models may hinder their scalability in OOD settings. Key factors, such as the time series embedding approach, attention mechanism, positional encoding, and forecasting strategy, all impact scalability to varying degrees. Among these factors, the embedding approach is particularly crucial, as it directly determines the input patterns fed into the model backbone and influences the operations the model can learn.
Theoretical Framework for Scaling Laws. Our scaling laws, similar to most scaling laws, are empirical findings. We believe a theoretical understanding of the learning dynamics that form the laws can provide a more solid justification. A potential perspective is understanding the optimization process through dynamics modeling (Bordelon et al., 2024). Specifically, based on the dynamical mean field theory, we can derive the infinite limit statistical description of the scaling behavior of training factors. By analyzing the response function in the infinite limit, which gauges the correlation between training factors and test error, we can obtain insights into how the scaling laws manifest.
Correlation Estimation for Metrics. Time series forecasting involves a diverse set of metrics, each measuring prediction error from a unique perspective. However, due to the lack of clear and definitive correlations between NLL loss and these metrics, it’s difficult to understand model prediction behavior directly from NLL loss. A potential perspective is to develop a prediction framework that bridges these metrics by empirically analyzing their relationships and dependencies (Isik et al., 2024). Such a framework would provide insights into how different metrics reflect model performance under varying conditions, enabling a deeper understanding of prediction behavior.